# Quantification of Interdependent Emotion Dynamics in Online Interactions

### v3

@@ Introduction
Using this framework, we analyze the joint dynamics of six basic emotions in YouTube live chats. We define the system of interest as the collective set of viewers participating in the live chat. Under this group-level framework, emotional expressions triggered by prior messages within the chat are considered endogenous, while influences originating outside the chat, primarily from the video content, are treated as exogenous. This distinction allows us to study how social interaction shapes group-level emotion dynamics during livestreams. Concretely, we classify emotions originating from video content as exogenous (exo) inputs and those expressed in the chat as endogenous (endo) responses driven by peer interaction. By estimating the parameters of a multivariate Hawkes model across 397 videos totaling more than 780 hours of content, we show that user emotional expressions are influenced by endogenous interactions roughly four times more than by exogenous video content. This suggests that in online environments, emotions are not simply reactions to external stimuli but are actively sustained and amplified through social feedback.
Our analysis further reveals a strong asymmetry in emotional contagion. Positive emotions, particularly joy, are about three times more likely to trigger additional joy than negative emotions are to trigger other negative expressions. However, this higher contagiousness is counterbalanced by slower temporal decay in negative emotions, which propagate over time windows nearly twice as long as those of positive ones. We also identify notable cross-excitation effects, especially cases where negative emotions trigger positive responses, a pattern reflective of antisocial dynamics such as trolling, where users derive gratification from provoking negative reactions [45, 46, 47].
Our study provides practical implications for the design of social networks and livestream platforms. First, increasing user exposure to joyful content can diffuse positivity across social communities. Second, active communication channels, such as live chats or bots, can much more effectively engage users emotionally compared to passive ones, such as video advertisements.
@@ Methods
We examine emotion dynamics in YouTube live videos with replay functionality across 27 topics ranging from comedy and documentary to sports and travel (see Appendix A for the complete list). We first compile the list of topic keywords and collect completed live video IDs for each topic keyword using the keyword search function via the YouTube API in Python. 111 Code and data are available at https://github.com/ivylyis/Quantification-of-the-Self-Excited-Emotion-Dynamics-in-Online-Interactions. We initially extract both the video transcripts and live chat messages associated with each video. This process yields two distinct lists: one of time-stamped chat messages and another of time-stamped video transcripts (subtitles). Notably, YouTube Live Chat offers two display modes: an unfiltered chronological stream (“Live Chat”) and a filtered version (“Top Chat”) that removes spam and inappropriate messages 222 https://support.google.com/youtube/answer/15268877?hl=en-GB . We use the “Live Chat” version in our data collection process, ensuring that all messages are collected without algorithmic filtering. While viewers can choose different display modes when watching the video, we note that YouTube’s filtering is not personalized and applies uniformly across all viewers. This stands in contrast to the personalized recommendation algorithms commonly used on other social media platforms, which can introduce more substantial distortions in the content users are exposed to.
@@
[…]
Multiple emotions can arise simultaneously in a single sentence [39, 41]. We thus assign to each text a vector of length 6 indicating the non-exclusive presence of each of the six emotions. This amounts to a multilabel classification problem. For training data, we rely on the SemEval-2018 dataset containing 6,838 Tweets with non-exclusive emotion labels across 11 emotions [53]. We group semantically related emotions into six core categories by mapping anticipation, optimism, love, and trust to joy; pessimism to sadness; and retaining sadness, anger, disgust, fear, and surprise in their original categories. Subsequently, we fine-tune a Roberta transformer model [54]. The transformer takes as input a text and gives as output the probabilities for each of the 6 basic emotions via 6 independent sigmoid activations. We retain emojis in the text input, as they serve as an important channel for emotional expression in online communication. We use the binary cross-entropy loss function for training. During prediction, we convert each sigmoid output to a binary indicator at the 0.5 cutoff. This allows us to assign to each live chat message and video subscript a subset of the emotions $\mathcal{E}$. We provide examples of live chat messages and the corresponding assigned emotion labels in Appendix A.
For every video and emotion analyzed, we determine the average count of live chat messages per minute. To exclude outliers or inconsistencies in live chat data, only those videos where this average falls between the $20^{\text{th}}$- and $80^{\text{th}}$-quantiles of all videos are included. This process results in a compiled set of 92,412 live chat messages that span 397 videos, totaling 780 hours. The median video length is 107 minutes and per video, the median number of live chat messages is 204. When considering all videos, the median number of live chat messages per minute is as follows: 1.63 for joy, 0.17 for disgust, 0.16 for anger, 0.11 for sadness, and for both fear and surprise, 0.03. Additional details are provided in Appendix A.
@@
Emotional contagion, the phenomenon in which emotions spread from one individual to another, is naturally captured by the self-exciting nature of the Hawkes process [55]. In our context, emotions are modeled as events whose statistical occurrence follows a point process. Because we have more than just one emotion and the emotions (presumably) cross-excite each other, we represent the dynamics of $|\mathcal{E}|=6$ distinct emotions using a multivariate Hawkes process [65]. It is our goal to understand to what extent a given emotion $f\in\mathcal{E}$ triggers another emotion $e\in\mathcal{E}$ either exogenously from the video to the live chat or endogenously from live chats to live chats. Here, the adjectives exogenous and endogenous refer to the point of view of live chat: If emotion $f$ appears in the video and triggers an emotion $e$ in the live chat, we call it exogenous (exo). In contrast, if emotion $f$ also appears in the live chat and subsequently triggers emotion $e$ in the live chat, we call it endogenous (endo). We assume that emotions in the video and in the live chat can trigger emotions in the live chat, but that emotions in the live chat cannot influence emotions in the video. We demonstrate in Appendix E that our results remain qualitatively consistent both with and without the subset of videos for which this assumption does not necessarily hold.
@@
$\scriptstyle\lambda^{e}(t)=\overbrace{\mu^{e}_{0}}^{\text{exo base rate}}+\overbrace{\sum_{f\in\mathcal{E}}\nu^{e,f}S^{f}(t)}^{\text{exo video influence}\leavevmode\nobreak\ \mu_{1}^{e}(t)}+\overbrace{\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\phi^{e,f}\left(t-t_{j}^{f}\right).}^{\text{endo chat influence}}$ (1)
@@
In summary, the intensity $\lambda^{e}(t)$ from Equation (1) is parametrized by a set of $142|\mathcal{E}|+2=$ 14 parameters: the exogeneous rate $\mu^{e}_{0}$, the decay time $\gamma^{e}$, as well as $|\mathcal{E}|$ parameters for $\nu^{e,f}$ and $\alpha^{e,f}$, respectively. We estimate these parameters by maximizing their log-likelihood functions (see below). Since each emotion has its own intensity $\lambda^{e}$, our model has a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters. However, the estimations of the different $\lambda^{e}$ are decoupled from each other, so that we can estimate each set of $12$ parameters separately. The fitted parameters are visualized in Figure 2 and discussed in the Results section.
@@
Previous studies have shown that assuming the exogenous influence $\mu^{e}(t)$ to be constant can lead to false attribution to endogenous effects [66, 60] (see Appendix B for additional confirmation). The YouTube live videos provide us with the unique opportunity to observe a time-varying influence of video content on user discussions. We use the video transcript to capture the time-varying emotional content of the video. For example, it is reasonable to assume that a particularly sad scene in the video induces sad emotions in the live chat. In addition, cross-influence can be expected; for example, a sad scene in the video induces angry emotions in the live chat (Figure 1). In particular, this is to be distinguished from a sad scene in the video triggering a sad message in the live chat, which in turn triggers either a sad or angry message. To capture such cross-influence from the videos, we parameterize the exogenous intensity of emotion $e$ as
$\mu^{e}(t)=\mu^{e}_{0}+\mu^{e}_{1}(t)\equiv\mu^{e}_{0}+\sum_{f\in\mathcal{E}}\leavevmode\nobreak\ \nu^{e,f}\leavevmode\nobreak\ S^{f}(t),$ (2)
where $\mu^{e}_{0}$ is a time-invariant, unobserved baseline intensity of spontaneous expression of emotion $e$ in the absence of influence from video or prior live chat messages. By contrast, we denote by $S^{f}(t)$ the time-varying intensity of emotion $f$ in the video, and $\nu^{e,f}$ the cross-influence from emotion $f$ in the video to emotion $e$ in the live chat. The term $\nu^{e,f}\leavevmode\nobreak\ S^{f}(t)$ is thus the time-varying intensity at which emotions of type $f$ in the video triggers emotions of type $e$ in the live chat.
@@
To fix $\mu$ and $\sigma$, we assume that the intensity of each transcript peaks $2$ seconds after appearance and that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance. In other words, we assume that the maximum of the log-normal function lies at $2$ seconds, $\exp(\mu-\sigma^{2})=2$, and that the median is equal to $10$ seconds, $\exp(\mu)=10$. Numerically solving for $\mu$ and $\sigma$ yields $2.3$ and $1.3$ respectively in units of minutes. To ensure that our results are not strongly dependent on this choice of $\mu$ and $\sigma$, we have checked that our results remain qualitatively similar for different values and functional shapes (see Appendix E). Alternatively, these parameters could be directly fitted from the data, a task which we leave for future research.
@@
$\log\left(L^{e}\right)$ $=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}^{e}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\lambda^{e}(s)}$ (4a)
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}^{e}))-\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\mu^{e}(s)}-$
$\quad\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {\phi^{e,f}(s)}.$ (4b)
@@
where $M^{f}_{1}\equiv\int\limits_{0}^{T}\leavevmode\nobreak\ \mathrm{d}s\leavevmode\nobreak\ {S^{f}(s)}$ is a constant that we calculate numerically. A detailed derivation of Equation (5) is found in Appendix C.
@@
Note that Equation (5) represents the log-likelihood for a single video, $\log L^{e}=\log L^{e}_{k}$ where $k$ enumerates the videos in our data sample. Thus, we aggregate $\log L^{e}_{k}$ across all videos and estimate the model parameters by maximizing $\sum_{k}\log L^{e}_{k}$. The total number of parameters is $142|\mathcal{E}|+2=$ 14 per emotion. This amounts to a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters to estimate for $|\mathcal{E}|=6$ emotions. However, the parameters for each emotion can be estimated independently because there is no interdependence between the parameters between different types of emotions. This gives rise to six independent fits of $14$ parameters each. To verify that the obtained fits are valid, we first test our model with synthetically generated data and obtain reliable performance (see Appendix D). We also refer to Appendix E for additional robustness checks with respect to different parameterizations and subsets of emotions. To mitigate potential biases, we estimate our model parameters using bootstrapped samples. We bootstrap our data $100$ times, each time sampling only 60% of all available videos. This allows us to estimate the parameters as averages across bootstrapped samples, and we use standard deviations as error bars. The fitted parameters are shown in Figure 2. These parameters are interpreted in the Results Section and discussed more qualitatively in the Discussion Section.
@@ Results
Figure 3: (a) For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time. Plot (a) shows the distribution of these ratios across videos. We notice that emotions are predominantly triggered endogenously. (b) Same as plot (a) but for the ratio of spontaneous (video-influenced) intensity to total exogenous intensity.
@@
Similarly, we can calculate the ratio of purely exogenous spontaneous emotions relative to video-induced emotions, that is $R^{e}_{0}=\mu^{e}_{0}/\left(\mu^{e}_{0}+\mu^{e}_{1}(t)\right)$ and its complement $R^{e}_{1}=1-R^{e}_{0}.$ Figure 3 (right) shows the distribution of $\langle R^{e}_{0}\rangle$ versus $\langle R^{e}_{1}\rangle$ over the set of all videos. The rate of chats induced by videos is typically one-fourth the rate of spontaneous chats. In summary, this highlights that live chat discussions elicit emotional expressions more effectively through social dynamics compared to passive video content consumption. In Appendix E, we further support these findings with additional robustness checks, varying different aspects of our study design, modeling selected subsets of the data sample and emotion categories. Below, we further show that there is some heterogeneity across video categories, which, however, does not affect the main conclusions of our study.
@@
Our main analysis examines general emotional and behavioral patterns among online users by compiling 92,412 YouTube live chat messages across 27 topics. The susceptibility of emotion dynamics to video influence and peer interactions is potentially subject to the type of video and the audience community. Here, we conduct sub-sample analyses on representative video types to identify heterogeneity in the evolution of online emotion dynamics. We focus on 4 video categories with keywords “politics”, “sports”, “live”, and “podcast”, with sample sizes 3,043, 11,172, 4,156, and 5,836 live chat messages, respectively. We report the model parameter estimates and relative ratios of different sources of emotion intensity in Appendix E.
Across different video categories, we observe several consistent patterns that align with our main findings: joy remains the most contagious emotion and exhibits the highest spontaneous intensity. Emotion dynamics are largely dominated by endogenous intensities except for the emotions of “surprise” and “fear”, which are shown to be more responsive to external stimuli, particularly for sport events and podcast shows. We also identify distinct patterns associated with specific video types and audience characteristics. The cross-triggering effects of anger and disgust are particularly potent for political videos, which contribute to moral outrage commonly observed in the political context. Unlike our main result (Figure 3), the content of political videos has a comparatively larger influence on the emotion dynamics of viewers, consistent with their provocative nature. In podcast and live-themed shows, surprise demonstrates a particularly long-lasting direct effect. Moreover, surprising content in podcast videos elicits joyful reactions. These findings validate our methodology by demonstrating its ability to disentangle the distinct channels through which video content and audience characteristics shape emotional dynamics.
In Appendix E, we show that our results remain robust when using alternative parameterizations of video influence, varying data preprocessing steps, excluding videos that primarily feature interactions between the audience and the video content, or modeling a reduced set of emotions.

### v4

# Social feedback amplifies emotional language in online video live chats

@@ Introduction
Using this framework, we analyze the joint dynamics of six basic emotions in YouTube live chats. We define the system of interest as the collective set of viewers participating in the live chat. Under this group-level framework, emotional expressions triggered by prior messages within the chat are considered endogenous, while influences originating outside the chat, primarily from the video content, are treated as exogenous. This distinction allows us to study how social interaction shapes group-level emotion dynamics during livestreams.
Concretely, we classify emotions originating from video content as exogenous (exo) inputs and those expressed in the chat as endogenous (endo) responses driven by peer interaction. Building on prior evidence of emotional contagion in digital environments, this study examines how emotional dynamics in live discussions arise from the interplay between external video content and social feedback. We address the extent to which emotional expressions are driven by exogenous video stimuli versus endogenous peer interactions. We further quantify the self- and cross-contagion effects as well as the temporal interdependencies of six different emotion types. These aims guide our empirical modeling of emotion dynamics using a multivariate Hawkes process framework applied to large-scale YouTube live chat data.
@@ Methods
We examine emotion dynamics in YouTube live videos with replay functionality across 27 topics ranging from comedy and documentary to sports and travel (see Supplementary Note 1 for the complete list). We first compile the list of topic keywords and collect completed live video IDs for each topic keyword using the keyword search function via the YouTube API in Python. We initially extract both the video transcripts and live chat messages associated with each video. This process yields two distinct lists: one of time-stamped chat messages and another of time-stamped video transcripts (subtitles). Notably, YouTube Live Chat offers two display modes: an unfiltered chronological stream (Live Chat) and a filtered version (Top Chat) that removes spam and inappropriate messages. We use the Live Chat version in our data collection process, ensuring that all messages are collected without algorithmic filtering. While viewers can choose different display modes when watching the video, we note that YouTube’s filtering is not personalized and applies uniformly across all viewers. This stands in contrast to the personalized recommendation algorithms commonly used on other social media platforms, which can introduce more substantial distortions in the content users are exposed to.
@@
We consider expressions of emotions through text. Text-based emotion measurement has been shown to correlate with self-reported emotions, establishing it as a reliable and valuable method for capturing emotional expressions [45, 46]. We delineate emotions according to the commonly used 6 basic emotions model [47]. These fundamental emotions are $\mathcal{E}\equiv\{\text{joy, surprise, anger, disgust, fear, sadness}\}$. The discrete emotion model remains a favored method within affective computing research [48]. Notably, the classification of emotions is effective for emotion analysis in textual content [49]. Moreover, research indicates that discrete emotion modeling often surpasses dimensional models in elucidating the sharing dynamics of online content, thereby more accurately capturing users’ emotional displays in digital settings [24].
Multiple emotions can arise simultaneously in a single sentence [39, 41]. We thus assign to each text a vector of length 6 indicating the non-exclusive presence of each of the six emotions. This amounts to a multilabel classification problem. For training data, we rely on the SemEval-2018 dataset containing 6,838 Tweets with non-exclusive emotion labels across 11 emotions [50]. We group semantically related emotions into six core categories by mapping anticipation, optimism, love, and trust to joy; pessimism to sadness; and retaining sadness, anger, disgust, fear, and surprise in their original categories. Subsequently, we fine-tune a Roberta transformer model [51]. The transformer takes as input a text and gives as output the probabilities for each of the 6 basic emotions via 6 independent sigmoid activations. We retain emojis in the text input, as they serve as an important channel for emotional expression in online communication. We use the binary cross-entropy loss function for training. During prediction, we convert each sigmoid output to a binary indicator at the 0.5 cutoff. This allows us to assign to each live chat message and video subscript a subset of the emotions $\mathcal{E}$. We provide examples of live chat messages and the corresponding assigned emotion labels in Supplementary Note 1.
For every video and emotion analyzed, we determine the average count of live chat messages per minute. To exclude outliers or inconsistencies in live chat data, only those videos where this average falls between the $20^{\text{th}}$- and $80^{\text{th}}$-quantiles of all videos are included. This process results in a compiled set of 92,412 live chat messages that span 397 videos, totaling 780 hours. The median video length is 107 minutes and per video, the median number of live chat messages is 204. When considering all videos, the median number of live chat messages per minute is as follows: 1.63 for joy, 0.17 for disgust, 0.16 for anger, 0.11 for sadness, and for both fear and surprise, 0.03. Additional details are provided in Supplementary Note 1.
@@
Emotional contagion, the phenomenon in which emotions spread from one individual to another, is naturally captured by the self-exciting nature of the Hawkes process [52]. In our context, emotions are modeled as events whose statistical occurrence follows a point process. Because we have more than just one emotion and the emotions (presumably) cross-excite each other, we represent the dynamics of $|\mathcal{E}|=6$ distinct emotions using a multivariate Hawkes process [62]. It is our goal to understand to what extent a given emotion $f\in\mathcal{E}$ triggers another emotion $e\in\mathcal{E}$ either exogenously from the video to the live chat or endogenously from live chats to live chats. Here, the adjectives exogenous and endogenous refer to the point of view of live chat: If emotion $f$ appears in the video and triggers an emotion $e$ in the live chat, we call it exogenous (exo). In contrast, if emotion $f$ also appears in the live chat and subsequently triggers emotion $e$ in the live chat, we call it endogenous (endo). We assume that emotions in the video and in the live chat can trigger emotions in the live chat, but that emotions in the live chat cannot influence emotions in the video. We demonstrate in Supplementary Note 5 that our results remain qualitatively consistent both with and without the subset of videos for which this assumption does not necessarily hold.
@@
$\scriptstyle\lambda^{e}(t)=\overbrace{\mu^{e}_{0}}^{\text{exo base rate}}+\overbrace{\sum_{f\in\mathcal{E}}\nu^{e,f}S^{f}(t)}^{\text{exo video influence}\penalty 10000\ \mu_{1}^{e}(t)}+\overbrace{\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}<t}\phi^{e,f}\left(t-t_{j}^{f}\right).}^{\text{endo chat influence}}$ (1)
@@
In summary, the intensity $\lambda^{e}(t)$ from Equation (1) is parametrized by a set of $2|\mathcal{E}|+2=14$ parameters: the exogeneous rate $\mu^{e}_{0}$, the decay time $\gamma^{e}$, as well as $|\mathcal{E}|$ parameters for $\nu^{e,f}$ and $\alpha^{e,f}$, respectively. We estimate these parameters by maximizing their log-likelihood functions (see below). Since each emotion has its own intensity $\lambda^{e}$, our model has a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters. However, the estimations of the different $\lambda^{e}$ are decoupled from each other, so that we can estimate each set of $12$ parameters separately. The fitted parameters are visualized in Figure 2 and discussed in the Results section.
We finally note that, in our analysis, we treat the aggregate of all viewers in the live chat as the endogenous system. At this level of description, any emotional comment triggered by prior comments in the chat is modeled as endogenous, while the influence of video content and a small baseline rate are treated as exogenous. This group-level framing is appropriate given our empirical setting, where repeated messages from the same viewer occur much more sparsely than messages exchanged across different viewers. If sufficient data were available, one could estimate a multivariate Hawkes process with $6N$ dimensions, where each of the $N$ viewers can experience six basic emotions. This would allow a precise separation between intra-individual self-excitation (diagonal terms of the kernel) and inter-individual contagion (off-diagonal terms). However, in such a specification, the number of parameters scale like $N^{2}$, which is computationally intensive and not feasible given our current dataset, where the average viewer posts only about nine messages per video. An intermediate approach would be to assume homogeneity across users and fit a single six-dimensional Hawkes process to all data, pooling observations across individuals. This would make it possible to quantify the average influence of other viewers and the video on a given individual, while keeping the parameter space manageable. Exploring this alternative model formulation represents a promising avenue for future research.
@@
Previous studies have shown that assuming the exogenous influence $\mu^{e}(t)$ to be constant can lead to false attribution to endogenous effects [63, 57] (see Supplementary Note 2 for additional confirmation). The YouTube live videos provide us with the unique opportunity to observe a time-varying influence of video content on user discussions. We use the video transcript to capture the time-varying emotional content of the video. For example, it is reasonable to assume that a particularly sad scene in the video induces sad emotions in the live chat. In addition, cross-influence can be expected; for example, a sad scene in the video induces angry emotions in the live chat (Figure 1). In particular, this is to be distinguished from a sad scene in the video triggering a sad message in the live chat, which in turn triggers either a sad or angry message. To capture such cross-influence from the videos, we parameterize the exogenous intensity of emotion $e$ as
$\mu^{e}(t)=\mu^{e}_{0}+\mu^{e}_{1}(t)\equiv\mu^{e}_{0}+\sum_{f\in\mathcal{E}}\penalty 10000\ \nu^{e,f}\penalty 10000\ S^{f}(t),$ (2)
where $\mu^{e}_{0}$ is a time-invariant, unobserved baseline intensity of spontaneous expression of emotion $e$ in the absence of influence from video or prior live chat messages. By contrast, we denote by $S^{f}(t)$ the time-varying intensity of emotion $f$ in the video, and $\nu^{e,f}$ the cross-influence from emotion $f$ in the video to emotion $e$ in the live chat. The term $\nu^{e,f}\penalty 10000\ S^{f}(t)$ is thus the time-varying intensity at which emotions of type $f$ in the video triggers emotions of type $e$ in the live chat.
@@
To fix $\mu$ and $\sigma$, we assume that the intensity of each transcript peaks $2$ seconds after appearance and that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance. In other words, we assume that the maximum of the log-normal function lies at $2$ seconds, $\exp(\mu-\sigma^{2})=2$, and that the median is equal to $10$ seconds, $\exp(\mu)=10$. Numerically solving for $\mu$ and $\sigma$ yields $2.3$ and $1.3$ respectively in units of minutes. To ensure that our results are not strongly dependent on this choice of $\mu$ and $\sigma$, we have checked that our results remain qualitatively similar for different values and functional shapes (see Supplementary Note 5). Alternatively, these parameters could be directly fitted from the data, a task which we leave for future research.
@@
$\log\left(L^{e}\right)$ $=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}^{e}))-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\lambda^{e}(s)}$ (4a)
$=\sum_{i=1}^{N^{e}}\log(\lambda^{e}(t_{i}^{e}))-\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\mu^{e}(s)}$-
$\quad\sum_{f\in\mathcal{E}}\sum_{t_{j}^{f}}\int\limits_{t_{j}^{f}}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {\phi^{e,f}(s)}.$ (4b)
@@
where $M^{f}_{1}\equiv\int\limits_{0}^{T}\penalty 10000\ \mathrm{d}s\penalty 10000\ {S^{f}(s)}$ is a constant that we calculate numerically. A detailed derivation of Equation (5) is found in Supplementary Note 3.
@@
Note that Equation (5) represents the log-likelihood for a single video, $\log L^{e}=\log L^{e}_{k}$ where $k$ enumerates the videos in our data sample. Thus, we aggregate $\log L^{e}_{k}$ across all videos and estimate the model parameters by maximizing $\sum_{k}\log L^{e}_{k}$. The total number of parameters is $2|\mathcal{E}|+2=14$ per emotion. This amounts to a total of $|\mathcal{E}|\left(2|\mathcal{E}|+2\right)=84$ parameters to estimate for $|\mathcal{E}|=6$ emotions. However, the parameters for each emotion can be estimated independently because there is no interdependence between the parameters between different types of emotions. This gives rise to six independent fits of $14$ parameters each. To verify that the obtained fits are valid, we first test our model with synthetically generated data and obtain reliable performance (see Supplementary Note 4). We also refer to Supplementary Note 5 for additional robustness checks with respect to different parameterizations and subsets of emotions. To mitigate potential biases, we estimate our model parameters using bootstrapped samples. We bootstrap our data $100$ times, each time sampling only 60% of all available videos. This allows us to estimate the parameters as averages across bootstrapped samples, and we use standard deviations as error bars. The fitted parameters are shown in Figure 2. These parameters are interpreted in the Results Section and discussed more qualitatively in the Discussion Section.
@@ Results
Figure 3: (a) For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time. Plot (a) shows the distribution of these ratios across $n=397$ videos. We notice that emotions are predominantly triggered endogenously. (b) Same as plot (a) but for the ratio of spontaneous (video-influenced) intensity to total exogenous intensity.
@@
Similarly, we can calculate the ratio of purely exogenous spontaneous emotions relative to video-induced emotions, that is $R^{e}_{0}=\mu^{e}_{0}/\left(\mu^{e}_{0}+\mu^{e}_{1}(t)\right)$ and its complement $R^{e}_{1}=1-R^{e}_{0}.$ Figure 3 (right) shows the distribution of $\langle R^{e}_{0}\rangle$ versus $\langle R^{e}_{1}\rangle$ over the set of all videos. The rate of chats induced by videos is typically one-fourth the rate of spontaneous chats. In summary, this highlights that live chat discussions elicit emotional expressions more effectively through social dynamics compared to passive video content consumption. In Supplementary Note 5, we further support these findings with additional robustness checks, varying different aspects of our study design, modeling selected subsets of the data sample and emotion categories. Below, we further show that there is some heterogeneity across video categories, which, however, does not affect the main conclusions of our study.
@@
Our main analysis examines general emotional and behavioral patterns among online users by compiling 92,412 YouTube live chat messages across 27 topics. The susceptibility of emotion dynamics to video influence and peer interactions is potentially subject to the type of video and the audience community. Here, we conduct sub-sample analyses on representative video types to identify heterogeneity in the evolution of online emotion dynamics. We focus on 4 video categories with keywords: politics, sports, live, and podcast, with sample sizes 3,043, 11,172, 4,156, and 5,836 live chat messages, respectively. We report the model parameter estimates and relative ratios of different sources of emotion intensity in Supplementary Note 5.
Across different video categories, we observe several consistent patterns that align with our main findings: joy remains the most contagious emotion and exhibits the highest spontaneous intensity. Emotion dynamics are largely dominated by endogenous intensities except for the emotions of surprise and fear, which are shown to be more responsive to external stimuli, particularly for sport events and podcast shows. We also identify distinct patterns associated with specific video types and audience characteristics. The cross-triggering effects of anger and disgust are particularly potent for political videos, which contribute to moral outrage commonly observed in the political context. Unlike our main result (Figure 3), the content of political videos has a comparatively larger influence on the emotion dynamics of viewers, consistent with their provocative nature. In podcast and live-themed shows, surprise demonstrates a particularly long-lasting direct effect. Moreover, surprising content in podcast videos elicits joyful reactions. These findings validate our methodology by demonstrating its ability to disentangle the distinct channels through which video content and audience characteristics shape emotional dynamics.
In Supplementary Note 5, we show that our results remain robust when using alternative parameterizations of video influence, varying data preprocessing steps, excluding videos that primarily feature interactions between the audience and the video content, or modeling a reduced set of emotions. In addition, we explicitly compare results between the full dataset and a sub-sample that excludes videos with highly interactive livestreamers. Although numerical parameter estimates differ slightly between these two sets, two-sample t-tests reveal no statistically significant differences, indicating that the distributions of estimates are indistinguishable. This suggests that feedback from the livestreamer itself plays a negligible role in shaping the overall emotional dynamics observed in our study.